Creative Computing sells a tablet computer called the Protab. The $780 sales price of a Protab Package includes the following:

One Protab computer.

A 6-month limited warranty. This warranty guarantees that Creative will cover any costs that arise due to repairs or replacements associated with defective products for up to six months.

A coupon to purchase a Creative Probook e-book reader for $200, a price that represents a 50% discount from the regular Probook price of $400. It is expected that 20% of the discount coupons will be utilized.

A coupon to purchase a one-year extended warranty for $50. Customers can buy the extended warranty for $50 at other times as well. Creative estimates that 40% of customers will purchase an extended warranty.

Creative does not sell the Protab without the limited warranty, option to purchase a Probook, and the option to purchase an extended warranty, but estimates that if it did so, a Protab alone would sell for $760.

Questions

1. How many performance obligations are included in a Protab Package? Explain your answer.

2. List the performance obligations in the Protab Package in the following table, and complete it to allocate the transaction price of 100,000 Protab Packages to the performance obligations in the contract. Creative Computing sells a tablet computer called

3.Prepare a journal entry to record sales of 100,000 Protab Packages (ignore any sales of extended warranties)
Answer #1
Concepts and reason

Performance obligation: It is defined as a promise to perform an activity related to goods or services. Whenever the two parties have signed the agreement for a consideration, then the parties involved in the agreement are bound to perform the activities for which the parties have been entered into the agreement.

Fundamentals

Service obligation: It is defined as the agreement between the two parties or more to perform a service for which the parties have entered into the agreement. The person who will provide the service is a service provider. The service provider is bound to provide the service required for the agreement. The party who will receive the service is service receiver. The service receiver is bound to make payment for the services received.

Limited warranty: It is a type of warranty in which a certain portion of products or certain kind of defects is considered under warranty. The limited warranty has certain conditions also. At many times, it only covers the cost of the materials and parts used while repairing the product. It does not cover the cost of labor services used to repair the product.

Separate performance obligation: The activity which is not bound to be purchased with the product or another service. It has a separate value. It can be sold separately. The performance obligation is the activity to perform the service. This service belongs to the product but can be sold separately with a specific value.

Service Revenue: It is the revenue that is generated by providing or performing services for the customer. Revenue will be recognized in the financial book when it is earned by the entity. The receipt of service revenue is not essential for recognizing the revenue.

Unearned Revenue: It is the outcome of accrual basis of accounting. When an income which is received during a specified period but it does not belong to that year or it is not earned during that year, then that revenue is treated as unearned revenue. It is considered as a liability of the organization until it is earned.

(1)

There are two performance obligations included in the Product P, which are given below:

Delivering the product P after selling is a performance obligation. 6 months limited warranty is not a separate performance obligation in which the obligation to perform the services is not separated from the product sold.

The coupon to purchase the C ProBook e-book reader is a separate performance obligation that will be utilized if the customer exercises the option.

The coupon to purchase the extended warranty for one year is not a performance obligation as no benefit, such as reduced price has been offered. The extended warranty in any case will cost $50, whether the customer buys it along with the protab or afterwards.

(2)

Prepare the table showing performance obligations, the selling price of the performance obligations and total transaction price of the performance obligations as given below:

Performance obligation
Stand alone selling
price of the
performance
obligation for
100,000 units
Percentage of the sum
of the

Working notes:

Compute performance obligation standalone’s sales price of the sale of Product P without limited warranty as given below:

Standalonesellingprice=Sellingpriceperunit×Unitssold=$760×100,000units=$76,000,000\begin{array}{c}\\{\rm{Standalone selling price}} = {\rm{Selling price per unit}} \times {\rm{Units sold}}\\\\ = \$ 760 \times 100,000{\rm{units}}\\\\ = \$ 76,000,000\\\end{array}

Compute performance obligation standalone sales price of the coupon to purchase the ProBook as given below:

Standalonesellingprice=Sellingpriceperunit×Unitssold=$200×100,000×20%=$200×20,000=$4,000,000\begin{array}{c}\\{\rm{Standalone selling price}} = {\rm{Selling price per unit}} \times {\rm{Units sold}}\\\\ = \$ 200 \times 100,000 \times 20\% \\\\ = \$ 200 \times 20,000\\\\ = \$ 4,000,000\\\end{array}

Compute performance obligation standalone sales price of the coupon to purchase the extended warranty as given below:

Standalonesellingprice=StandaloneSellingpriceperunit×Unitssold=($50$50)×100,000×40%=0\begin{array}{c}\\{\rm{Standalone selling price}} = {\rm{Standalone Selling price per unit}} \times {\rm{Units sold}}\\\\ = \left( {\$ 50 - \$ 50} \right) \times 100,000 \times 40\% \\\\ = 0\\\end{array}

Compute total of performance obligations for selling 100,000 units as given below:

Totalsalesof100,000units=SalesofProductP+OptiontopurchaseProbook+Optiontopurchaseextendedwarranty=$76,000,000+$4,000,000+0=$80,000,000\begin{array}{c}\\{\rm{Total sales of 100,000units}} = {\rm{Sales of Product P}} + {\rm{Option to purchase Probook}}\\\\ + {\rm{Option to purchase extended warranty}}\\\\ = \$ 76,000,000 + \$ 4,000,000 + 0\\\\ = \$ 80,000,000\\\end{array}

Compute the percentage of the sale of product P with the sum of the performance obligation stand-alone selling prices as given below:

%ofsaleofProductP=SellingpriceofsingleperformanceobligationSellingpriceoftotalperformanceobligations×100=$76,000,000$80,000,000×100=95%\begin{array}{c}\\{\rm{\% of sale of Product P }} = \frac{{{\rm{Selling price of single performance obligation}}}}{{{\rm{Selling price of total performance obligations}}}} \times 100\\\\ = \frac{{\$ 76,000,000}}{{\$ 80,000,000}} \times 100\\\\ = 95\% \\\end{array}

Compute the percentage of the coupon to purchase the ProBook with the sum of the performance obligation stand-alone selling prices as given below:

%ofcoupontopurchasetheprobook=SellingpriceofsingleperformanceobligationSellingpriceoftotalperformanceobligations×100=$4,000,000$80,000,000×100=5%\begin{array}{c}\\{\rm{\% of coupon to purchase the probook}} = \frac{\begin{array}{l}\\{\rm{Selling price of single }}\\\\{\rm{performance obligation}}\\\end{array}}{\begin{array}{l}\\{\rm{Selling price of total }}\\\\{\rm{performance obligations}}\\\end{array}} \times 100\\\\ = \frac{{\$ 4,000,000}}{{\$ 80,000,000}} \times 100\\\\ = 5\% \\\end{array}

Compute the total transaction price of the sale of Product P according to the percentage of allocation as given below:

Totaltransactionprice=Percentageoftransaction×Transactionprice×Unitssold=95%×$780×100,000units=95%×$78,000,000=$74,100,000\begin{array}{c}\\{\rm{Total transaction price}} = {\rm{Percentage of transaction}} \times {\rm{Transaction price}} \times {\rm{Units sold}}\\\\ = 95\% \times \$ 780 \times 100,000{\rm{units}}\\\\ = 95\% \times \$ 78,000,000\\\\ = \$ 74,100,000\\\end{array}

Compute the total transaction price of the coupon to purchase the ProBook according to the percentage of allocation as given below:

Totaltransactionprice=Percentageoftransaction×Transactionprice×Unitssold=5%×$780×100,000units=5%×$78,000,000=$3,900,000\begin{array}{c}\\{\rm{Total transaction price}} = {\rm{Percentage of transaction}} \times {\rm{Transaction price}} \times {\rm{Units sold}}\\\\ = 5\% \times \$ 780 \times 100,000{\rm{units}}\\\\ = 5\% \times \$ 78,000,000\\\\ = \$ 3,900,000\\\end{array}

Compute total transaction price of all the performance obligations as given below:

Totaltransactionprice=TransactionpriceofSaleofproductP+Transactionpriceofcoupontopurchaseprobook=$74,100,000+$3,900,000=$78,000,000\begin{array}{c}\\{\rm{Total transaction price}} = {\rm{Transaction price of Sale of product P}}\\\\ + {\rm{Transaction price of coupon to purchase probook}}\\\\{\rm{ = \$ 74,100,000}} + \$ 3,900,000\\\\ = \$ 78,000,000\\\end{array}

(3)

Journal entry to record the sale of Protab is given below:

Journal
Date Account title and Explanation
Cash
| Sales Revenue
Deferred Revenue Discount Option
(To record the sale of prota

Ans: Part 1

There are two performance obligations which forms the package of Protab.

The transaction value of protab without limited warranty is $74,100,000 and coupons to purchase the e-book reader is $3,900,000.

Part 3

Journal
Date Account title and Explanation
Cash
| Sales Revenue
Deferred Revenue Discount Option
(To record the sale of prota
Answer #2

Been looking for this answer for hours, finally. The only right answer here. Thank you so much frr


answered by: anonymous